# Singleton algorithms for the Constraint Satisfaction Problem

### v1

@@ 1.4 Structure of the paper
The paper is organized as follows. In Section 2 we give necessary definitions. Section 3 introduces classic universal algorithms and their combinations, defines palette block symmetric functions, and formulates criteria for these algorithms to solve $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$. In Section 4 we discuss the limitations and weaknesses of the classic universal algorithms such as BLP, AIP, and BLP+AIP, and demonstrate why the singleton versions are more powerful. Additionally, we show that the template coming from the dihedral group $\mathbb{D}_{4}$ cannot be solved by the singleton version of $\mathrm{BLP}+\mathrm{AIP}$. In Section 5 we prove the relational characterization of the dihedral group $\mathbb{D}_{4}$ and show that it does not admit palette block symmetric polymorphisms. In Section 6 we prove the characterization of the singleton versions of the algorithms in terms of existence of palette block symmetric polymorphisms. Finally, in Section 7 we prove that every finite template on a domain of size at most 7 that admits a WNU polymorphism also admits a palette block symmetric polymorphism, and therefore can be solved by the singleton version of $\mathrm{BLP}+\mathrm{AIP}$.
@@ 2 Preliminaries
A function $f\colon A^{n}\to B$ is called a polymorphism from a $\sigma$-structure $\mathbb{A}$ to $\sigma$-structure $\mathbb{B}$ if, for each symbol $R\in\sigma$ and tuples $\mathbf{a}_{1},\dots,\mathbf{a}_{n}\in R^{\mathbb{A}}$, we have $f(\mathbf{a}_{1},\dots,\mathbf{a}_{n}):=(f(\mathbf{a}_{1}^{1},\dots,\mathbf{a}_{n}^{1}),\dots,f(\mathbf{a}_{1}^{\operatorname{ar}(R)},\dots,\mathbf{a}_{n}^{\operatorname{ar}(R)}))\in R^{\mathbb{B}}$. By $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ we denote the set of all polymorphisms from $\mathbb{A}$ to $\mathbb{B}$. When $\mathbb{A}=\mathbb{B}$, we write simply $\operatorname{Pol}(\mathbb{A})$.
@@
For a set $N$, we denote by $\mathrm{id}_{N}\colon N\to N$ the identity map. For a relation $R\subseteq A^{n}$ and $S\subseteq[n]$ by $\operatorname{pr}_{S}(R)$ we denote the projection of $R$ onto the coordinates indexed by $S$.
@@ 3.1 Arc-consistency algorithm
The algorithm $\mathrm{ArcCons}$ returns “Yes” on a CSP instance $\mathcal{I}$ if and only if there exists a mapping $D^{\prime}$ that assigns a nonempty set $D_{x}^{\prime}\subseteq D_{x}$ to every variable $x$ such that for any constraint $R(x_{1},\dots,x_{s})$ of $\mathcal{I}$ and any $j\in[s]$ we have $\operatorname{pr}_{j}(R\cap(D_{x_{1}}^{\prime}\times\dots\times D_{x_{s}}^{\prime}))=D_{x_{j}}^{\prime}$.
The power of the arc-consistency algorithm has been completely characterized for both $\operatorname{CSP}(\mathbb{A})$ and $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$. An function $f\colon A^{n}\to B$ is called totally symmetric if $f(a_{1},\dots,a_{n})=f(b_{1},\dots,b_{n})$ whenever $\{a_{1},\dots,a_{n}\}=\{b_{1},\dots,b_{n}\}$.
@@
Let $(\mathbb{A},\mathbb{B})$ be a PCSP template. Then $\mathrm{ArcCons}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains totally symmetric functions of all arities.
@@ 3.2 Linear relaxations
The next algorithms come from the idea of applying the algorithms for solving linear equations (such as Gaussian elimination) to arbitrary CSPs over finite domains. Suppose $\mathcal{I}$ is a CSP instance given by the triple $\langle X,D,\mathbf{C}\rangle$. By $\mathcal{I}^{L}$ we denote the following system of linear equations (the meaning of + is not specified, as it is treated as a formal expression):
@@ 3.3 Combinations of the algorithms
Similarly, by $\mathrm{ArcCons}$$\mathfrak{A}$ we denote the algorithm that first runs the arc-consistency algorithm, returns “No” if the latter returns “No”, and otherwise applies $\mathfrak{A}$ to the reduced instance and returns its answer.
@@
3.4 Singleton algorithms and palette block functions
@@
It is not generally true that $\mathrm{SinglAlg}$ is weaker than $\mathrm{CSingl}\mbox{{Alg}}$, since the algorithm $\mathrm{SinglAlg}$, when fixing a variable $x$ to a value, may reduce all constraints containing $x$, while $\mathrm{CSingl}\mbox{{Alg}}$ always reduces only one constraint. Nevertheless, if $\mbox{{Alg}}\in\{\mathrm{ArcCons},\mathrm{BLP},\mathrm{ArcCons}+\mathfrak{A},\mathrm{BLP}+\mathfrak{A}\}$, then $\mathrm{SinglAlg}$ is clearly weaker than $\mathrm{CSingl}\mbox{{Alg}}$.
@@
Similarly to block functions, we define palette block functions, where we require the block properties only on palette tuples. Formally, a function $f\colon A^{k_{1}+\dots+k_{n}}\to A$ is called $\mathcal{P}$-$(k_{1},\dots,k_{n})$-palette block function, and the $i$-th block of it is (totally) symmetric/(weakly) alternating if (1) and (2) imply (3):
@@
A function is $\mathcal{P}$-$(k_{1},\dots,k_{n})$-palette block (totally) symmetric/(weakly) alternating if it is a $\mathcal{P}$-$(k_{1},\dots,k_{n})$-palette block function whose blocks are (totally) symmetric/(weakly) alternating.
The power of singleton algorithms and their combinations can be characterized using palette block functions. Below we present a characterization for the algorithms $\mathrm{Singl}\mathrm{ArcCons}$ and $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$, which are proved in Section 6.3 and 6.4, respectively.
@@
    $\mathrm{CSingl}\mathrm{ArcCons}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
@@
    $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has an $(\overline{m_{1}},\dots,\overline{m_{n}})$-palette block totally symmetric function for every $n,m_{1},\dots,m_{n}\in\mathbb{N}$;
@@
    for every $N\in\mathbb{N}$ there exist $n,m_{1},\dots,m_{n}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has an $(\overline{m_{1}},\dots,\overline{m_{n}})$-palette block totally symmetric function.
@@
    $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette block symmetric function for every $\ell,n\in\mathbb{N}$;
@@
    for every $N\in\mathbb{N}$ there exist $n\geq N$, $k_{1},\dots,k_{n}\in\mathbb{N}$, and $m_{1,1},\dots,m_{1,k_{1}},\dots,m_{n,1},\dots,m_{n,k_{n}}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\overline{m_{1,1},\dots,m_{1,k_{1}}},\overline{m_{2,1},\dots,m_{2,k_{2}}},\dots,\overline{m_{n,1},\dots,m_{n,k_{n}}})$-palette block symmetric function.
To explain the idea behind palette block symmetric functions and Theorem 3.6, let us prove the following claim.
@@
Suppose $(\mathbb{A},\mathbb{B})$ is a PCSP template, and for every $L\in\mathbb{N}$ there exist $n,\ell_{1},\dots,\ell_{n}\geq L$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette block symmetric function. Then $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$.
@@
It is sufficient to prove that if $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ returns “Yes” on an instance $\mathcal{I}$, then $\mathcal{I}$ has a solution. For every variable $x$ and every element $a$ from the obtained domain of $x$ let $\mathcal{I}_{x}^{a}$ be the instance obtained in the algorithm SinglAlg by reducing the domain of $x$ to $\{a\}$. Let $\Omega$ be the set of all such instances. Since $\mathrm{BLP}+\mathrm{AIP}$ returns “Yes” on each of these instances, the fact (f3) from Section 3.3 implies that there exists a large number $L$ such that $L>|\Omega|$ and each instance $\mathcal{I}_{i}^{a}$ has a $d$-solution for any $d\geq L$. Choose an $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette block symmetric function $f\in\operatorname{Pol}(\mathbb{A},\mathbb{B})$, where $n,\ell_{1},\dots,\ell_{n}\geq L$.
@@
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ are similar Taylor algebras, $|A_{i}|<8$ for every $i\in[s]$. Then $\mathbf{A}_{1}\times\dots\times\mathbf{A}_{s}$ has a $(\underbrace{p^{m},\dots,p^{m}}_{n})$-palette block symmetric term operation for every prime $p>8$ and every $m,n\in\mathbb{N}$.
@@
Most other combinations of the algorithms can be characterized similarly (see Table 1). Since their proofs are almost identical to those for $\mathrm{Singl}\mathrm{ArcCons}$ and $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$, we omit them. We can also characterize the algorithm $\mathrm{Singl}\mathfrak{A}\wedge\mathfrak{B}$ for $\mathfrak{A},\mathfrak{B}\in\{\mathrm{ArcCons},\mathrm{BLP},\mathrm{AIP}\}$, but this requires a different notion and is less interesting, since the algorithms are simply run independently. For example, $\mathrm{Singl}\mathrm{BLP}\wedge\mathrm{AIP}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if for all $m,\ell,n\in\mathbb{N}$, $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains a $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},\ell)$-block function such that the first $n$ blocks are symmetric on $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},\ell)$-palette tuples and the last block is always alternating.

### v2

@@ 1.4 Structure of the paper
The paper is organized as follows. In Section 2 we give necessary definitions. Section 3 introduces classic universal algorithms and their combinations, defines palette symmetric functions, and formulates criteria for these algorithms to solve $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$. In Section 4 we discuss the limitations and weaknesses of the classic universal algorithms such as BLP, AIP, and BLP+AIP, and demonstrate why the singleton versions are more powerful. Additionally, we show that the template coming from the dihedral group $\mathbf{D}_{4}$ cannot be solved by the singleton version of $\mathrm{BLP}+\mathrm{AIP}$. In Section 5 we prove the characterization of the singleton versions of the algorithms in terms of existence of palette symmetric polymorphisms. Section 6 provides concrete palette symmetric polymorphisms for temporal relational structures and shows that all tractable temporal CSPs can be solved (after sampling) by $\mathrm{Singl}\mathrm{AIP}$ or $\mathrm{CSingl}\mathrm{AC}$. In Section 7 we prove the relational characterization of the dihedral group $\mathbf{D}_{4}$ and show that it does not admit palette symmetric polymorphisms. Finally, in Section 8 we prove that every finite template on a domain of size at most 7 that admits a WNU polymorphism also admits a palette symmetric polymorphism, and therefore can be solved by the singleton version of $\mathrm{BLP}+\mathrm{AIP}$.
@@ 2 Preliminaries
A function $f\colon A^{n}\to B$ is called a polymorphism from a $\sigma$-structure $\mathbb{A}$ to a $\sigma$-structure $\mathbb{B}$ if, for each symbol $R\in\sigma$ and tuples $\mathbf{a}_{1},\dots,\mathbf{a}_{n}\in R^{\mathbb{A}}$, we have $f(\mathbf{a}_{1},\dots,\mathbf{a}_{n}):=(f(\mathbf{a}_{1}^{1},\dots,\mathbf{a}_{n}^{1}),\dots,f(\mathbf{a}_{1}^{\operatorname{ar}(R)},\dots,\mathbf{a}_{n}^{\operatorname{ar}(R)}))\in R^{\mathbb{B}}$. By $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ we denote the set of all polymorphisms from $\mathbb{A}$ to $\mathbb{B}$. When $\mathbb{A}=\mathbb{B}$, we write simply $\operatorname{Pol}(\mathbb{A})$.
@@
For a set of operations $F\subseteq\mathcal{O}_{A}$ by $\operatorname{Clo}(F)$ we denote the minimal clone containing $F$. For a set $N$, we denote by $\mathrm{id}_{N}\colon N\to N$ the identity map. For a relation $R\subseteq A^{n}$ and $S\subseteq[n]$ by $\operatorname{pr}_{S}(R)$ we denote the projection of $R$ onto the coordinates indexed by $S$.
@@ 3.1 Arc-consistency algorithm
The algorithm $\mathrm{AC}$ returns “Yes” on a CSP instance $\mathcal{I}$ if and only if there exists a mapping $D^{\prime}$ that assigns a nonempty set $D_{x}^{\prime}\subseteq D_{x}$ to every variable $x$ such that for any constraint $R(x_{1},\dots,x_{s})$ of $\mathcal{I}$ and any $j\in[s]$ we have $\operatorname{pr}_{j}(R\cap(D_{x_{1}}^{\prime}\times\dots\times D_{x_{s}}^{\prime}))=D_{x_{j}}^{\prime}$.
The power of the arc-consistency algorithm has been completely characterized for both $\operatorname{CSP}(\mathbb{A})$ and $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$. A function $f\colon A^{n}\to B$ is called totally symmetric if $f(a_{1},\dots,a_{n})=f(b_{1},\dots,b_{n})$ whenever $\{a_{1},\dots,a_{n}\}=\{b_{1},\dots,b_{n}\}$.
@@
Let $(\mathbb{A},\mathbb{B})$ be a PCSP template. Then $\mathrm{AC}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains totally symmetric functions of all arities.
One direction of this theorem can be proved as follows. Suppose $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains totally symmetric polymorphisms of all arities. First, we choose a sequence of totally symmetric polymorphisms $f_{1},f_{2},f_{n},\dots,$ such that each $f_{n}$ is of arity $n$, and identifying two variables in $f_{n}$ yields $f_{n-1}$. Then, for a set $S=\{a_{1},\dots,a_{s}\}\subseteq A$, we define $f(S):=f_{s}(a_{1},\dots,a_{s})$. If the algorithm $\mathrm{AC}$ returns “Yes” on a CSP instance $\mathcal{I}$, then it assigns a nonempty set $D_{x}^{\prime}$ to every variable $x$. We claim that a solution to $\mathcal{I}^{\mathbb{B}}$ can be defined by $s(x):=f(D_{x}^{\prime})$ for every variable $x$. Let $R(x_{1},\dots,x_{k})$ be a constraint $\mathcal{I}$. Consider $R^{\mathbb{A}}\cap(D_{x_{1}}^{\prime}\times\dots\times D_{x_{k}}^{\prime})=\{\mathbf{a}_{1},\dots,\mathbf{a}_{m}\}$. Since $f_{m}$ is a polymorphism, we have $f_{m}(\mathbf{a}_{1},\dots,\mathbf{a}_{m})\in R^{\mathbb{B}}$. By the definition of $f$, we obtain $f_{m}(\mathbf{a}_{1},\dots,\mathbf{a}_{m})=(s(x_{1}),\dots,s(x_{k}))$. Therefore, $(s(x_{1}),\dots,s(x_{k}))\in R^{\mathbb{B}}$. Thus, $s$ is a solution of $\mathcal{I}$ in $\mathbb{B}$, and $\mathrm{AC}$ solves the instance correctly.
@@ 3.2 Linear relaxations
The next algorithms arise from the idea of applying algorithms for solving linear equations (such as Gaussian elimination) to arbitrary CSPs over finite domains. Suppose $\mathcal{I}$ is a CSP instance given by the triple $\langle X,D,\mathbf{C}\rangle$. By $\mathcal{I}^{L}$ we denote the following system of linear equations (the meaning of + is not specified, as it is treated as a formal expression):
@@ 3.3 Combinations of the algorithms
Similarly, by $\mathrm{AC}$$\mathfrak{A}$ we denote the algorithm that first runs the arc-consistency algorithm, returns “No” if the latter returns “No”, and otherwise applies $\mathfrak{A}$ to the reduced instance and returns its answer.
@@
3.4 Singleton algorithms and palette functions
@@
It is not generally true that $\mathrm{SinglAlg}$ is weaker than $\mathrm{CSingl}\mbox{{Alg}}$, since the algorithm $\mathrm{SinglAlg}$, when fixing a variable $x$ to a value, may reduce all constraints containing $x$, while $\mathrm{CSingl}\mbox{{Alg}}$ always reduces only one constraint. Nevertheless, if $\mbox{{Alg}}\in\{\mathrm{AC},\mathrm{BLP},\mathrm{AC}+\mathfrak{A},\mathrm{BLP}+\mathfrak{A}\}$, then $\mathrm{SinglAlg}$ is clearly weaker than $\mathrm{CSingl}\mbox{{Alg}}$.
@@
Similarly to block functions, we define palette functions, where we require the block properties only on palette tuples. Formally, a function $f\colon A^{k_{1}+\dots+k_{n}}\to A$ is called $\mathcal{P}$-$(k_{1},\dots,k_{n})$-palette function, and the $i$-th block of it is (totally) symmetric/(weakly) alternating if (1) and (2) imply (3):
@@
A function is $\mathcal{P}$-$(k_{1},\dots,k_{n})$-palette (totally) symmetric/(weakly) alternating if it is a $\mathcal{P}$-$(k_{1},\dots,k_{n})$-palette function whose blocks are (totally) symmetric/(weakly) alternating.
The power of singleton algorithms and their combinations can be characterized using palette functions. Below we present a characterization for the algorithms $\mathrm{Singl}\mathrm{AC}$ and $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$, which are proved in Section 5.3 and 5.4, respectively.
@@
    $\mathrm{CSingl}\mathrm{AC}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
@@
    $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has an $(\overline{m_{1}},\dots,\overline{m_{n}})$-palette totally symmetric function for every $n,m_{1},\dots,m_{n}\in\mathbb{N}$;
@@
    for every $N\in\mathbb{N}$ there exist $n,m_{1},\dots,m_{n}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has an $(\overline{m_{1}},\dots,\overline{m_{n}})$-palette totally symmetric function.
@@
    $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette symmetric function for every $\ell,n\in\mathbb{N}$;
@@
    for every $N\in\mathbb{N}$ there exist $n\geq N$, $k_{1},\dots,k_{n}\in\mathbb{N}$, and $m_{1,1},\dots,m_{1,k_{1}},\dots,m_{n,1},\dots,m_{n,k_{n}}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\overline{m_{1,1},\dots,m_{1,k_{1}}},\overline{m_{2,1},\dots,m_{2,k_{2}}},\dots,\overline{m_{n,1},\dots,m_{n,k_{n}}})$-palette symmetric function.
To explain the idea behind palette symmetric functions and Theorem 3.6, let us prove the following claim.
@@
Suppose $(\mathbb{A},\mathbb{B})$ is a PCSP template, and for every $L\in\mathbb{N}$ there exist $n,\ell_{1},\dots,\ell_{n}\geq L$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette symmetric function. Then $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$.
@@
It is sufficient to prove that if $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ returns “Yes” on an instance $\mathcal{I}$, then $\mathcal{I}$ has a solution. For every variable $x$ and every element $a$ from the obtained domain of $x$ let $\mathcal{I}_{x}^{a}$ be the instance obtained in the algorithm SinglAlg by reducing the domain of $x$ to $\{a\}$. Let $\Omega$ be the set of all such instances. Since $\mathrm{BLP}+\mathrm{AIP}$ returns “Yes” on each of these instances, the fact (f3) from Section 3.3 implies that there exists a large number $L$ such that $L>|\Omega|$ and each instance $\mathcal{I}_{i}^{a}$ has a $d$-solution for any $d\geq L$. Choose an $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette symmetric function $f\in\operatorname{Pol}(\mathbb{A},\mathbb{B})$, where $n,\ell_{1},\dots,\ell_{n}\geq L$.
@@
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ are similar algebras admitting WNU term operations, $|A_{i}|<8$ for every $i\in[s]$. Then $\mathbf{A}_{1}\times\dots\times\mathbf{A}_{s}$ has a $(\underbrace{p^{m},\dots,p^{m}}_{n})$-palette symmetric term operation for every prime $p>8$ and every $m,n\in\mathbb{N}$.
@@
Most other combinations of the algorithms can be characterized similarly (see Table 1). Since their proofs are almost identical to those for $\mathrm{Singl}\mathrm{AC}$ and $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$, we omit them. We can also characterize the algorithm $\mathrm{Singl}\mathfrak{A}\wedge\mathfrak{B}$ for $\mathfrak{A},\mathfrak{B}\in\{\mathrm{AC},\mathrm{BLP},\mathrm{AIP}\}$, but this requires a different notion and is less interesting, since the algorithms are simply run independently. For example, $\mathrm{Singl}\mathrm{BLP}\wedge\mathrm{AIP}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if for all $m,\ell,n\in\mathbb{N}$, $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains a $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},\ell)$-block function such that the first $n$ blocks are symmetric on $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},\ell)$-palette tuples and the last block is always alternating.